# A review of Bayesian sensor-based estimation and uncertainty quantification of aerodynamic flows

### v2

@@ 3.2 Nonlinear operators
Alternatively, if we assume that the distributions remain reasonably compact, then we can assume that they remain Gaussian. In that case, we could linearize each operator about the most recent mean state and use the standard Kalman filter (30) and (32). In this approach, called the extended Kalman filter, the matrix operators $F_{k}$ and $H_{k}$ are replaced by Jacobians of the nonlinear operators evaluated at the respective means. Hemati et al. [17] used such an approach for estimating the parameters of line vortices that represented the wake of a lead aircraft in a formation, from pressure sensor data along the span of a trailing aircraft. These sensor measurements were interpreted with an observation model built from the Bernoulli equation and a lifting-line model of the trailing aircraft’s own wake distribution. The extended Kalman filter may perform poorly when the operators are highly nonlinear, since a single Jacobian evaluated at the mean is insufficient for characterizing the uncertainty propagation. Instead, the unscented Kalman filter [18] represents each state distribution with deterministically-chosen “sigma points”, at which we evaluate the operators and then construct the necessary covariances to apply the Kalman steps.
For systems of large state dimension $n$, as we frequently encounter in fluid dynamics, neither the extended nor the unscented Kalman filter is practical, since they both require explicit construction of covariance matrices that are $n\times n$. Some researchers have addressed this by working in a reduced-order basis for the flow field. Fukumori and Malanotte-Rizzoli [19] proposed this approach in the context of oceanographic flows, and it was subsequently adapted by Suzuki [20] to estimate shear flows by assimilating velocimetry data into a coarse direct numerical simulation (DNS). We will have more to say about reduced-order bases in Section 4.
Instead of computing these covariances explicitly, we could pursue something more like the particle filter, using an ensemble of states to approximate the distributions. But to avoid the computational drawbacks of the particle filter, we will assume that these distributions remain Gaussian, enabling us to draw upon some aspects of the Kalman filter. This is the foundation of the ensemble Kalman filter (EnKF), first proposed by Evensen [21] in the context of weather forecasting. The basic algorithm is as follows: For an ensemble of states $\{\mathsf{x}^{i}\}_{i=1}^{M}$ drawn from some prior distribution, the forecast step is just a propagation of each ensemble member $i=1,\ldots,M$ through the forecast operator,
@@
where the construction of the Kalman gain $K$ can be done in a few ways, most of which make use of the fact that the exact Kalman gain is $K=\Sigma_{\mathsf{x}\mathsf{y}}\Sigma_{\mathsf{y}\mathsf{y}}^{-1}$, as shown in (54). In particular, in the basic stochastic EnKF (sEnKF) [21], the Kalman gain is simply computed using ensemble approximations of these covariances, $K=\hat{\Sigma}_{\mathsf{x}\mathsf{y}}\hat{\Sigma}_{\mathsf{y}\mathsf{y}}^{-1}$, where the approximations are given by (61) and (62), using the ensemble that emerges from the forecast step. The advantage of the EnKF for high-dimensional problems (with ensembles smaller than the state dimension, $M<n$) is that the computational costs of the forecast and analysis steps scale much better than for particle filters, since they simply involve $M$ evaluations of the forecast and observation operators, can be implemented using only existing codes for these operators, and require less storage than extended and unscented Kalman filters.
@@
The sEnKF unfortunately exhibits a few undesirable behaviors for small- to modest-sized ensembles, where the computational advantages of the EnKF can best be realized. Much of this degraded performance can be attributed to two issues associated with a finite-sized ensemble: underestimation of the random process error in the forecast and spurious correlations in the ensemble covariances. The first of these issues is a common outcome of a deterministic forecast model (in which the process noise in (34) is not naturally present) and causes a phenomenon called covariance collapse: the ensemble members all converge to the same value after several time steps, so the covariance matrices (60) and (61) (and thus, $K$) tend toward zero, preventing further updates of the ensemble from future measurements. It has become common practice to address this with a simple trick called “covariance inflation”, which comes in two forms [3]: a multiplicative form, in which each ensemble member’s anomaly (i.e., its deviation from the mean) is multiplied by a factor $\beta$ just larger than 1; and an additive form, in which a random process noise $\eta^{i}\sim\mathcal{N}(0,\Sigma_{\eta})$ is added to each ensemble member. The choices for $\beta$ and $\Sigma_{\eta}$ need not be tied to anything physical; they simply ensure a minimum level of uncertainty so that the ensemble remains responsive to new measurements.
@@
The sEnKF with inflation has been used successfully in several flow estimation investigations. Colburn et al. [22] used pressure and skin friction measurements at the wall of an incompressible turbulence channel flow to estimate the mean velocity field in the channel. Da Silva and Colonius [23] used it to estimate the evolving velocity field of a pair of low Reynolds number flows: past a circular cylinder, using velocity sensors in the wake; and past a NACA 0009 airfoil at 30 degrees angle of attack, using surface pressure sensors. In both cases, they used direct numerical simulation of the incompressible Navier–Stokes equations and associated pressure Poisson equation to provide the forecast and observation operators. Truth measurements were provided by the same simulations, but with noise added to the observations. In the work of Darakananda et al. [24], surface pressure measurements were obtained (with noise added) from direct numerical simulation of the Navier–Stokes equations of a flat-plate airfoil at 20 and 60 degrees angle of attack and Reynolds number 500, in some cases subjected to large impulsive flow disturbances. They interpreted these measurements with an inviscid point vortex model of the same geometry. An ensemble of 50 members were propagated with this model via the sEnKF with inflation. In this work, the estimator’s knowledge of the flow disturbance is only obtained from the surface measurements, and these were shown to successfully update the vortices’ positions and strengths in the analysis step to mimic the disturbance’s convective influence.
@@
The relatively large ensemble size needed in [24] reflects the second issue with sEnKF, due to spurious correlations. This issue can be addressed in a few ways, depending on the origin. For example, the measurement noise $\varepsilon^{i}$ and state $\mathsf{x}^{i}$, though ideally uncorrelated, are not uncorrelated in a finite ensemble. A particular consequence is that the posterior and prior covariances do not obey the ideal relationship in (10), $\Sigma_{\mathsf{x}|\mathsf{y}^{\ast}}=(I_{n}-KH)\Sigma_{\mathsf{x}}$, that expresses the analysis step’s reduction of our uncertainty. Thus, many of the pernicious effects of noise–state correlation are mitigated with the ensemble transform Kalman filter (ETKF), developed by Bishop et al. [25], in which the Kalman update is re-derived to exactly reproduce the ideal relationship for any ensemble size. Le Provost and Eldredge [26] showed that the ETKF can significantly improve the performance of the sEnKF for aerodynamic flows similar to those considered in [24]. An example of those results is shown in Figure 6.
Figure 6: Top panels: Comparison at three instants of the true vorticity field from DNS of a Reynolds number 500 flow past a flat plate at 20 degrees angle of attack (upper row) and the mean vortex positions of an ensemble of vortex models that assimilate surface pressure from the DNS, using the ETKF (lower row). Lower panel: comparison of true normal force (black line) with the estimated force (orange, with 95 confidence interval). Adapted from [26].
@@
The spurious correlations among states and between states and observations are more complicated to address because they are necessarily mixed in with true physical correlations. The technique of “localization” [3] can often effectively address this by restricting covariance entries to those in which the state and the sensor are close in some geometric sense. However, many observations in fluid dynamics involve algebraic relationships over long distances, e.g., pressure and velocity, and localization can artificially truncate these relationships. Thus, we present a variant of the EnKF in the next section that provides an alternative approach to offset the challenges from spurious correlations.
@@ 3.3 The low-rank EnKF
Before we present the low-rank EnKF (or LREnKF) [12], it is important to note that the ensemble covariances (60)–(62) are limited to rank $M-1$, as discussed below those equations in B. Thus, if the number of sensors and the state dimension are both larger than the ensemble size, the state update (35) in the sEnKF—by virtue of (61)—is rank deficient, limited to a linear combination of the $M$ state anomalies $\mathsf{x}^{i}-\hat{\mu}_{\mathsf{x}}$ in the prior ensemble (where $\hat{\mu}_{\mathsf{x}}$ is the ensemble mean (58)). Without further treatment, the sEnKF weights this linear combination haphazardly, in a manner that can be corrupted by the aforementioned spurious correlations. It would be better if we could curate this linear combination to avoid these spurious influences.
How can we curate it? We saw in the case of the linear operator in Section 2 that the columns of $V$ and $U$ form respective bases for the state and measurement spaces that, by the magnitude of the singular values, partition the informative from the noninformative subspaces of these spaces; for a nonlinear observation operator in an ensemble, we can make use of the eigenvectors and corresponding eigenvalues of the Gramians (63) and (64) for this purpose, as we discussed in Section 2.4. It is important to stress that the “noninformative” subspace is not concretely defined and can include directions with eigenvalues that are relatively small compared to the dominant ones. In fact, the most pernicious aspects of the finite-sized ensemble typically lie in these smaller-eigenvalue directions, so we expect better behavior if we simply avoid them. This is the foundational idea of the LREnKF [12]. We truncate each basis at respective ranks $r_{\mathsf{x}}$ and $r_{\mathsf{y}}$, chosen so that the cumulative fraction of energy in the retained directions just exceeds some threshold, $\alpha$, i.e., $\sum_{j=1}^{r_{\mathsf{x}}}\lambda_{\mathsf{x},j}/\sum_{j=1}^{n}\lambda_{\mathsf{x},j}\geq\alpha$ and $\sum_{j=1}^{r_{\mathsf{y}}}\lambda_{\mathsf{y},j}/\sum_{j=1}^{d}\lambda_{\mathsf{y},j}\geq\alpha$, in which the eigenvalues are sorted in decreasing order. After choosing the reduced ranks, the Kalman gain is constructed via the spectrally-decomposed form (21), but now with the reduced bases $V_{r_{\mathsf{x}}}$ and $U_{r_{\mathsf{y}}}$:
@@
There are two complications of the LREnKF compared to the sEnKF. The LREnKF needs access to the Jacobian of the observation operator, ideally by analytical means or automatic differentiation, which is more intrusive than the basic sEnKF. Furthermore, one must compute eigenvectors and eigenvalues of the Gramian matrices, which are of size $\mathbb{R}^{n\times n}$ and $\mathbb{R}^{d\times d}$, respectively. For large-dimensional systems, the first of these can be computationally intensive. However, the method enables the use of very small ensembles (smaller than either $n$ or $d$) without exhibiting filter divergence, as occurs for the sEnKF in such ensemble sizes [12]. Thus, there is a trade-off in cost that makes these extra computations worthwhile.
@@ 3.4 Example: sequential estimation of point vortices
In Section 2.3.2 we discussed an example in which we used pressure measurements to infer the characteristics of one or more point vortices at a single instant. In that example, we assumed very little prior knowledge and used MCMC sampling to characterize the posterior probability. In the current example, we will also estimate point vortices from pressure measurements. However, now we will perform the estimation over an extended time interval as the flow evolves, taking in pressure measurements at a constant rate and using the LREnKF to assimilate these measurements into an ensemble that approximates the evolving probability. In this case, we will consider a true flow as depicted in Fig. 7, consisting of point vortices in a uniform flow of speed $U$ past a circular cylinder, and in which measurements are obtained from a uniformly-distributed array of pressure sensors on the cylinder surface. For the purposes of demonstration, we will assume the flow is inviscid, incompressible, and two-dimensional. We obviously recognize that this true flow is not representative of a realistic vortex encounter in a viscous flow. Our purpose here is merely to illuminate the general aspects of the observability of vortices.
As in the previous example, the state vector includes the positions and strengths of the vortices. The forecast operator comprises the time advancement of the positions of the point vortices by a forward Euler integration of the local velocity field for one time step; the observation is obtained from solution of the pressure Poisson equation. The velocity and pressure fields are obtained with a grid-based potential flow solver that uses an immersed boundary treatment of the body surface and lattice Green’s function for the solution of Poisson equation; the details are described in [27]. The grid spacing is $\Delta x/R_{c}=0.02$ and the time step size is $U\Delta t/R_{c}=0.02$. The Jacobian of the observation operator, $\nabla h$, is evaluated using automatic differentiation of the aforementioned numerical operators. True pressure measurements $\mathsf{y}^{\ast}_{k}$ are provided at every time step, and these are synthetically generated with the same forecast and observation operators as in the estimation, but starting from an initial state $\mathsf{x}_{0}$ unknown to the estimator and augmented with Gaussian measurement noise with standard deviation $\sigma_{\mathcal{E}}=10^{-3}$.

### v3

@@ 3.2 Nonlinear operators
Alternatively, if we assume that the distributions remain reasonably compact, then we can assume that they remain Gaussian. In that case, we could linearize each operator about the most recent mean state and use the standard Kalman filter (30) and (32). In this approach, called the extended Kalman filter, the matrix operators $F_{k}$ and $H_{k}$ are replaced by Jacobians of the nonlinear operators evaluated at the respective means. Hemati et al. [hemati2014wake] used such an approach for estimating the parameters of line vortices that represented the wake of a lead aircraft in a formation, from pressure sensor data along the span of a trailing aircraft. These sensor measurements were interpreted with an observation model built from the Bernoulli equation and a lifting-line model of the trailing aircraft’s own wake distribution. The extended Kalman filter may perform poorly when the operators are highly nonlinear, since a single Jacobian evaluated at the mean is insufficient for characterizing the uncertainty propagation. Instead, the unscented Kalman filter [wan2000unscented] represents each state distribution with deterministically-chosen “sigma points”, at which we evaluate the operators and then construct the necessary covariances to apply the Kalman steps.
For systems of large state dimension $n$, as we frequently encounter in fluid dynamics, neither the extended nor the unscented Kalman filter is practical, since they both require explicit construction of covariance matrices that are $n\times n$. Some researchers have addressed this by working in a reduced-order basis for the flow field. Fukumori and Malanotte-Rizzoli [fukumori1995approximate] proposed this approach in the context of oceanographic flows, and it was subsequently adapted by Suzuki [suzuki2012reduced] to estimate shear flows by assimilating velocimetry data into a coarse direct numerical simulation (DNS). We will have more to say about reduced-order bases in Section 4.
Instead of computing these covariances explicitly, we could pursue something more like the particle filter, using an ensemble of states to approximate the distributions. But to avoid the computational drawbacks of the particle filter, we will assume that these distributions remain Gaussian, enabling us to draw upon some aspects of the Kalman filter. This is the foundation of the ensemble Kalman filter (EnKF), first proposed by Evensen [evensen1994sequential] in the context of weather forecasting. The basic algorithm is as follows: For an ensemble of states $\{\mathsf{x}^{i}\}_{i=1}^{M}$ drawn from some prior distribution, the forecast step is just a propagation of each ensemble member $i=1,\ldots,M$ through the forecast operator,
@@
where the construction of the Kalman gain $K$ can be done in a few ways, most of which make use of the fact that the exact Kalman gain is $K=\Sigma_{\mathsf{x}\mathsf{y}}\Sigma_{\mathsf{y}\mathsf{y}}^{-1}$, as shown in (54). In particular, in the basic stochastic EnKF (sEnKF) [evensen1994sequential], the Kalman gain is simply computed using ensemble approximations of these covariances, $K=\hat{\Sigma}_{\mathsf{x}\mathsf{y}}\hat{\Sigma}_{\mathsf{y}\mathsf{y}}^{-1}$, where the approximations are given by (61) and (62), using the ensemble that emerges from the forecast step. The advantage of the EnKF for high-dimensional problems (with ensembles smaller than the state dimension, $M<n$) is that the computational costs of the forecast and analysis steps scale much better than for particle filters, since they simply involve $M$ evaluations of the forecast and observation operators, can be implemented using only existing codes for these operators, and require less storage than extended and unscented Kalman filters.
@@
The sEnKF unfortunately exhibits a few undesirable behaviors for small- to modest-sized ensembles, where the computational advantages of the EnKF can best be realized. Much of this degraded performance can be attributed to two issues associated with a finite-sized ensemble: underestimation of the random process error in the forecast and spurious correlations in the ensemble covariances. The first of these issues is a common outcome of a deterministic forecast model (in which the process noise in (34) is not naturally present) and causes a phenomenon called covariance collapse: the ensemble members all converge to the same value after several time steps, so the covariance matrices (60) and (61) (and thus, $K$) tend toward zero, preventing further updates of the ensemble from future measurements. It has become common practice to address this with a simple trick called “covariance inflation”, which comes in two forms [asch2016data]: a multiplicative form, in which each ensemble member’s anomaly (i.e., its deviation from the mean) is multiplied by a factor $\beta$ just larger than 1; and an additive form, in which a random process noise $\eta^{i}\sim\mathcal{N}(0,\Sigma_{\eta})$ is added to each ensemble member. The choices for $\beta$ and $\Sigma_{\eta}$ need not be tied to anything physical; they simply ensure a minimum level of uncertainty so that the ensemble remains responsive to new measurements.
@@
The sEnKF with inflation has been used successfully in several flow estimation investigations. Colburn et al. [colburn2011state] used pressure and skin friction measurements at the wall of an incompressible turbulent channel flow to estimate the mean velocity field in the channel. Da Silva and Colonius [da2018ensemble] used it to estimate the evolving velocity field of a pair of low Reynolds number flows: past a circular cylinder, using velocity sensors in the wake; and past a NACA 0009 airfoil at 30 degrees angle of attack, using surface pressure sensors. In both cases, they used direct numerical simulation of the incompressible Navier–Stokes equations and associated pressure Poisson equation to provide the forecast and observation operators. Truth measurements were provided by the same simulations, but with noise added to the observations. In the work of Darakananda et al. [darakjdeprf18], surface pressure measurements were obtained (with noise added) from direct numerical simulation (DNS) of the Navier–Stokes equations of a flat-plate airfoil at 20 and 60 degrees angle of attack and Reynolds number 500, in some cases subjected to large impulsive flow disturbances. They interpreted these measurements with an inviscid point vortex model of the same geometry. An ensemble of 50 members was propagated with this model via the sEnKF with inflation. In that work, the estimator’s knowledge of the flow disturbance was only obtained from the surface pressure measurements of the DNS, and these were shown to successfully update the vortices’ positions and strengths in the analysis step to mimic the disturbance’s convective influence.
@@
The relatively large ensemble size needed in [darakjdeprf18] reflects the second issue with sEnKF, due to spurious correlations. This issue can be addressed in a few ways, depending on the origin. For example, the measurement noise $\varepsilon^{i}$ and state $\mathsf{x}^{i}$, though ideally uncorrelated, are not uncorrelated in a finite ensemble. A particular consequence is that the posterior and prior covariances do not obey the ideal relationship in (10), $\Sigma_{\mathsf{x}|\mathsf{y}^{\ast}}=(I_{n}-KH)\Sigma_{\mathsf{x}}$, that expresses the analysis step’s reduction of our uncertainty. Thus, many of the pernicious effects of noise–state correlation are mitigated with the ensemble transform Kalman filter (ETKF), developed by Bishop et al. [bishop2001adaptive], in which the Kalman update is re-derived to exactly reproduce the ideal relationship for any ensemble size. Le Provost and Eldredge [le2021ensemble] showed that the ETKF can significantly improve the performance of the sEnKF for aerodynamic flows similar to those considered in [darakjdeprf18]. An example of those results is shown in Figure 6, which depicts, at three instants during the disturbance response, a comparison of the true vorticity from DNS with the ensemble mean vortex positions of the estimator. A comparison of estimated and true normal force is also shown.
Figure 6: Comparison of true flow (DNS, Re = 500) and point-vortex estimate, from DNS surface pressure. Adapted with permission from Le Provost and Eldredge (Phys. Rev. Fluids, 2021) [le2021ensemble], copyright 2021 American Physical Society.
@@
The spurious correlations among states and between states and observations are more complicated to address because they are necessarily mixed in with true physical correlations. The technique of “localization” [asch2016data] can often effectively address this by restricting covariance entries to those in which the state and the sensor are close in some geometric sense. However, many observations in fluid dynamics involve algebraic relationships over long distances, e.g., pressure and velocity, and localization can artificially truncate these relationships. Thus, we present a variant of the EnKF in the next section that provides an alternative approach to offset the challenges from spurious correlations.
@@ 3.3 The low-rank EnKF
Before we present the low-rank EnKF (or LREnKF) [le2022low], it is important to note that the ensemble covariances (60)–(62) are limited to rank M1M-1, as discussed below those equations in Appendix B. Thus, if the number of sensors and the state dimension are both larger than the ensemble size, the state update (35) in the sEnKF—by virtue of (61)—is rank deficient, limited to a linear combination of the $M$ state anomalies $\mathsf{x}^{i}-\hat{\mu}_{\mathsf{x}}$ in the prior ensemble (where $\hat{\mu}_{\mathsf{x}}$ is the ensemble mean (58)). Without further treatment, the sEnKF weights this linear combination haphazardly, in a manner that can be corrupted by the aforementioned spurious correlations. It would be better if we could curate this linear combination to avoid these spurious influences.
How can we curate it? We saw in the case of the linear operator in Section 2 that the columns of $V$ and $U$ form respective bases for the state and measurement spaces that, by the magnitude of the singular values, partition the informative from the noninformative subspaces of these spaces; for a nonlinear observation operator in an ensemble, we can make use of the eigenvectors and corresponding eigenvalues of the Gramians (63) and (64) for this purpose, as we discussed in Section 2.4. It is important to stress that the “noninformative” subspace is not concretely defined and can include directions with eigenvalues that are relatively small compared to the dominant ones. In fact, the most pernicious aspects of the finite-sized ensemble typically lie in these smaller-eigenvalue directions, so we expect better behavior if we simply avoid them. This is the foundational idea of the LREnKF [le2022low]. We truncate each basis at respective ranks $r_{\mathsf{x}}$ and $r_{\mathsf{y}}$, chosen so that the cumulative fraction of energy in the retained directions just exceeds some threshold, $\alpha$, i.e., $\sum_{j=1}^{r_{\mathsf{x}}}\lambda_{\mathsf{x},j}/\sum_{j=1}^{n}\lambda_{\mathsf{x},j}\geq\alpha$ and $\sum_{j=1}^{r_{\mathsf{y}}}\lambda_{\mathsf{y},j}/\sum_{j=1}^{d}\lambda_{\mathsf{y},j}\geq\alpha$, in which the eigenvalues are sorted in decreasing order. After choosing the reduced ranks, the Kalman gain is constructed via the spectrally-decomposed form (21), but now with the reduced bases $V_{r_{\mathsf{x}}}$ and $U_{r_{\mathsf{y}}}$:
@@
There are two complications of the LREnKF compared to the sEnKF. The LREnKF needs access to the Jacobian of the observation operator, ideally by analytical means or automatic differentiation, which is more intrusive than the basic sEnKF. Furthermore, one must compute eigenvectors and eigenvalues of the Gramian matrices, which are of size $\mathbb{R}^{n\times n}$ and $\mathbb{R}^{d\times d}$, respectively. For large-dimensional systems, the first of these can be computationally intensive. However, the method enables the use of very small ensembles (smaller than either $n$ or $d$) without exhibiting filter divergence, as occurs for the sEnKF in such ensemble sizes [le2022low]. Thus, there is a trade-off in cost that makes these extra computations worthwhile.
@@ 3.4 Example: sequential estimation of point vortices
In Section 2.3.2 we discussed an example in which we used pressure measurements to infer the characteristics of one or more point vortices at a single instant. In that example, we assumed very little prior knowledge and used MCMC sampling to characterize the posterior probability. In the current example, we will also estimate point vortices from pressure measurements. However, now we will perform the estimation over an extended time interval as the flow evolves, taking in pressure measurements at a constant rate and using the LREnKF to assimilate these measurements into an ensemble that approximates the evolving probability. In this case, we will consider a true flow as depicted in Fig. 7, consisting of point vortices self-advecting in a uniform flow of speed $U$ past a circular cylinder, and in which measurements are obtained from a uniformly-distributed array of pressure sensors on the cylinder surface. For the purposes of demonstration, we will assume the flow is inviscid, incompressible, and two-dimensional. We obviously recognize that this true flow is not representative of a realistic vortex encounter in a viscous flow. Our purpose here is merely to illuminate the general aspects of the observability of vortices.
As in the previous example, the state vector includes the positions and strengths of the vortices. The forecast operator comprises the time advancement of the positions of the point vortices by a forward Euler integration of the local velocity field for one time step; the observation is obtained from solution of the pressure Poisson equation. The velocity and pressure fields are obtained with a grid-based potential flow solver that uses an immersed boundary treatment of the body surface and lattice Green’s function for the solution of Poisson equation; the details are described in [beckers2022planar]. The grid spacing is $\Delta x/R_{c}=0.02$ and the time step size is $U\Delta t/R_{c}=0.02$. The Jacobian of the observation operator, $\nabla h$, is evaluated using automatic differentiation of the aforementioned numerical operators. True pressure measurements $\mathsf{y}^{\ast}_{k}$ are provided at every time step, and these are synthetically generated with the same forecast and observation operators as in the estimation, but starting from an initial state $\mathsf{x}_{0}$ unknown to the estimator and augmented with Gaussian measurement noise with standard deviation $\sigma_{\mathcal{E}}=10^{-3}$.